OpenStudy (anonymous):

Please help I am stuck Use part I of the Fundamental Theorem of Calculus to find the derivative of the integral from 4x to 5x of (u+5/u-1)du I will also show this as an equation below

OpenStudy (anonymous):

\[\int\limits_{4x}^{5x}\frac{u+5 }{ u-1 }du\]

OpenStudy (anonymous):

So far I know that I can separate this into two different integrals \[\int\limits_{0}^{5x}\frac{ u+5 }{ u-1}du -\int\limits_{0}^{4x}\frac{ u+5 }{ u-1 }\]